OpenStudy (anonymous):

What is x^2+x-6 factored out

OpenStudy (jhannybean):

what two numbers multiply to give you -1 but add to give you -6?

OpenStudy (daniellelovee):

(x+3)(x-2)

OpenStudy (anonymous):

x^2+x-6 =x^2+3x-2x-6 =x(x+3)-2(x+3) =(x+3)(x-2)